# DiCoFlex: Model-agnostic diverse counterfactuals with flexible control

### v1

@@ 1 Introduction
Counterfactual explanations (CFs) have become an integral part of explainable artificial intelligence (XAI) by providing human-interpretable insights into complex machine learning models [10]. CFs answer critical "what-if" questions by suggesting minimal and meaningful changes to input data that could alter the outcome of a predictive model [40]. Such explanations have valuable applications across sensitive domains, including finance, healthcare, and legal decisions, where understanding the predictions of the model and potential alternative outcomes is paramount [42, 9].
@@
According to Guidotti [10], an ideal counterfactual explanation should satisfy several key properties. First, it must demonstrate validity by successfully changing the model’s prediction. Second, it should maintain proximity by remaining as close as possible to the original input, minimizing the amount of change. Third, it should exhibit sparsity by modifying only the smallest possible set of features. Fourth, it must ensure plausibility by remaining within the realm of realistic, observed data. Finally, it should guarantee actionability by suggesting changes that are feasible for the stakeholder to implement.
In practice, however, these criteria often conflict. For instance, the smallest possible change may result in samples outside the realistic data distribution. To address this, Mothilal et al. [19] and Laugel et al. [17] argue that, rather than producing a single “best” counterfactual explanation, a diverse set of CFs should be generated. Presenting multiple realistic alternatives allows users to see different ways to achieve the desired outcome, explore trade-offs between minimal change and realism, and gain deeper insight into the model’s decision boundaries. This diversity, in turn, empowers stakeholders to make more informed, feasible decisions.
Current approaches to diverse counterfactual generation face significant limitations. Optimization-based methods like DiCE [19] and multi-objective frameworks [5] incorporate diversity directly into their cost functions but require solving complex, separate optimizations for each explanation. This results in high computational costs and latency, making them impractical for real-time applications. Meanwhile, generative approaches such as those by Pawelczyk et al. [25] and Panagiotou et al. [22] offer efficiency improvements but frequently produce redundant explanations, require access to model gradients, and lack flexibility for adapting to user-defined constraints without extensive retraining.
@@ 2 Related Work
Counterfactual explanations (CFs) have become integral to explainable AI by providing intuitive alternatives that elucidate model decisions [40]. Optimization-based methods formulate CF generation as minimizing distance between instances while ensuring prediction changes, using implementations including gradient-based approaches [40], integer programming [37], genetic algorithms [5], and satisfiability problems [15]. Instance-based methods like FACE [26] select examples yielding desired predictions from existing data, while CBCE [16] merges features from paired contrasting instances. Model-specific approaches such as decision tree methods [11, 35] leverage model structure by traversing alternative paths.
@@
Generating multiple diverse counterfactuals offers significant advantages over single explanations by providing alternative paths for recourse and better illustrating decision boundaries [17, 19]. DiCE [19] directly incorporates diversity into its loss function but requires separate optimization for each explanation, increasing computational costs. Other approaches achieve diversity through various strategies: DiVE [30] employs determinantal point processes, while methods like [36, 32, 4] partition the feature space. Multi-objective approaches like MOC [5] generate solutions representing different trade-offs between properties such as proximity and sparsity, while CARE [27] and OrdCE [14] incorporate user preferences into the generation process.
@@
Generative models have gained popularity for counterfactual generation due to their ability to produce more plausible explanations that remain on the data manifold. VAE-based approaches include ReViSE [13], C-CHVAE [25], CLUE [1], and CRUDS [8], which learn latent representations of the data distribution. GAN-based methods such as GeCo [33] and SCGAN [42] use adversarial training to balance validity and realism. Recent transformer-based approaches like TABCF [22] specifically target tabular data with mixed types. However, many of these approaches still require access to model gradients, produce redundant explanations, or lack flexibility in handling user-defined constraints.
@@
Practical counterfactual explanations must satisfy various constraints including sparsity (minimizing modified features) and actionability (ensuring feasible changes). CERTIFAI [34] employs genetic algorithms to enforce constraints, while MCCE [28] conditions on immutable features. Methods focused on actionable recourse, such as CSCF [20], account for downstream effects of feature modifications, while Russell [31] and Verma et al. [39] generate explanations representing different modification strategies. However, most existing approaches require retraining or reoptimization when constraints change, limiting their flexibility in real-world applications.
@@
Current counterfactual methods face substantial limitations for tabular data: optimization-based approaches require high computational costs for diverse explanations [19], and generative models often produce redundant explanations, require model gradients, or lack constraint flexibility [25]. Our proposed approach, DiCoFlex, addresses these limitations through a conditional generative framework that efficiently produces diverse counterfactuals in a single forward pass, enables real-time constraint customization, operates without requiring model access, and naturally handles mixed feature types.
@@ 3.2 Modeling predictive conditional distribution
Our approach aims at directly optimizing the conditional negative log-likelihood given by (4) with the data sampled from the distribution $q(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})$. As a consequence, it requires access to the parametrized density function. Several techniques, such as Kernel Density Estimation (KDE) and Gaussian Mixture Models (GMM), can be used to model the conditional density function $p_{\theta}(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})$. In this work, we propose using a conditional normalizing flow model [29] for this purpose. Unlike KDE or GMM, normalizing flows do not rely on a predefined parametric form of the density function and are well-suited for modeling high-dimensional data. Furthermore, in contrast to other generative models, such as diffusion models or GANs, normalizing flows allow for exact density evaluation via the change of variables formula and can be efficiently trained by minimizing the negative log-likelihood. The general formula that represents normalizing flows can be expressed as:
@@
where $f_{\theta}(\mathbf{x}^{\prime};\mathbf{x},y^{\prime})$ is an invertible transformation conditioned on $\mathbf{x}$ and $y^{\prime}$, $p_{Z}(f_{\theta}(\mathbf{x}^{\prime};\mathbf{x},y^{\prime}))$ is the base distribution, typically Gaussian and $J_{f_{\theta}}(\mathbf{x}^{\prime};\mathbf{x},y^{\prime})$ is Jacobian of the transformation. The biggest challenge in normalizing flows is the choice of the invertible function for which the determinant of Jacobian is easy to calculate. Several solutions have been proposed in the literature to address this issue with notable approaches, including NICE [6], RealNVP [7], and MAF [23]. Based on superior performance on benchmark dataset, we selected MAF as our normalizing flow model (see Appendix C).
3.3 Scoring training counterfactual explanations
The distribution $q(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})$ is essential to sample data to train the flow model. Since we do not have direct access to this distribution, we introduce a score function, which quantifies what "the adequate counterfactual" means. Following the basic postulates defined by [10], the counterfactual has to satisfy at least validity, proximity, and plausibility constraints. These requirements lead us to the following score function:
$s(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})=\log p(y^{\prime}|\mathbf{x}^{\prime})-\lambda_{1}d(\mathbf{x},\mathbf{x}^{\prime})+\lambda_{2}p_{\text{data}}(\mathbf{x}^{\prime}),$ (6)
where $p(y^{\prime}|\mathbf{x}^{\prime})$ is class probability returned by model $h(\cdot)$, $p_{\text{data}}(\mathbf{x}^{\prime})$ is value of density function for $\mathbf{x}^{\prime}$ and $\lambda_{1},\lambda_{2}$ are user-defined trade-off parameters. Making use of the above criterion, we can evaluate the quality of a counterfactual candidate.
There are several approaches for including the above score function to the learning process of normalizing flow modeling the distribution of counterfactual explanations. However, to keep our model as simple as possible, we resort to the straightforward technique of querying training examples, which are similar to the input $\mathbf{x}$ and satisfy $h(\mathbf{x}^{\prime})=y^{\prime}$. For this purpose, we use a typical KNN (K-nearest neighbors) heuristic, in which we sample $K$ data points labeled as $y^{\prime}$ by the classifier that are closest to the input $\mathbf{x}$. Since we are sampling training examples from class $y^{\prime}$, the validity and plausibility constraints hold, while the proximity is satisfied by using the KNN approach.
@@ 3.3 Scoring training counterfactual explanations
[…]
@@
One could ask whether a direct use of the distribution $\hat{q}(\mathbf{x}^{\prime}\mid\mathbf{x},y^{\prime},d)$ can be an alternative for generating counterfactual explanations, instead of learning a parametrized flow-based model ${p_{\theta}(\mathbf{x}^{\prime}\mid\mathbf{x},y^{\prime})}$. Observe that the use of $\hat{q}$ is inherently limited by the number of available training examples and lacks the ability to generalize to unseen instances. Additionally, the diversity of generated examples is constrained by the number of selected neighbors. Modeling a generative model provides smooth changes over input examples and allows for extrapolating counterfactual generation outside training data.
@@ 3.4 Sparsity on continuous features
$d_{p}(\mathbf{x}_{i}^{\prime},\mathbf{x})=||\mathbf{x}_{i}^{\prime}-\mathbf{x}||_{p}=\left(\sum_{j=1}^{D}|x^{\prime}_{j}-x_{j}|^{p}\right)^{1/p}.$ (8)
We can observe that the level of modified attributes can be controlled through an appropriate choice of the value of $p$. For instance, if $p=0.01$, neighbors with fewer modified attributes are favored, thereby enforcing sparsity. Making use of $p=2$ gives a standard Euclidean proximity measure. Since categorical attributes are commonly encoded by one-hot vectors, the above scheme affects mostly continuous features.
In practice, the sparsity level can be effectively controlled in the inference time by incorporating the parameter $p$ into the flow model $p_{\theta}(\mathbf{x}^{\prime}\mid\mathbf{x},y^{\prime},p)$ through additional conditioning. During training, values of $p$ are sampled and explicitly included in the conditioning of the model, allowing it to learn the relationship between $p$ and the desired sparsity in the generated outputs. As a result, during the inference stage, users can adjust the sparsity level of the modified attributes simply by selecting an appropriate value of $p$. This mechanism provides a flexible and interpretable way to regulate the extent of changes applied to the input $\mathbf{x}$, depending on specific needs or constraints. Note, however, that the applied $L_{p}$ norm does not freeze the continuous attributes but only encourages small changes on their portion, which can be seen as a type of soft constraint.
@@ 3.6 Training
Concluding, the training procedure for estimating $\theta$ is as follows. For each training iteration, $\mathbf{x}$ is selected from data examples and the target class $y^{\prime}$ is sampled for counterfactual explanation. Value $p$ that represents the sparsity level is also sampled from some predefined set $\mathcal{P}$, as well as the masking vector $\mathbf{m}$ from $\mathcal{M}$. Next, the set $N(\mathbf{x},y^{\prime},d_{p,\mathbf{m}},K)$ of $K$ nearest neighbors is calculated considering distance measure $d_{p,\mathbf{m}}$ given by (9). The counterfactual $\mathbf{x}^{\prime}$ is sampled from $\hat{q}(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime},d_{p,\mathbf{m}})$ given by (7). Finally, $\theta$ is updated in gradient-based procedure by optimizing (4), considering conditioning both on $p$ and m. Appendix D contains the details of the training algorithm.
@@ 4 Experiments
In this section, we conduct a series of experiments that evaluate the quality of counterfactual explanations generated by various methods. Our main objective is to understand how effectively these counterfactuals flip the predictions of a model while maintaining realism, requiring minimal changes to the input, and offering diversity among the generated alternatives.
@@ 4.1 Experimental Setup
We evaluate all methods on five well-established benchmark datasets commonly used in counterfactual explanation research: Adult [2] (income prediction), Bank Marketing [18] (customer response classification), Default [41] (credit card default risk prediction), Give Me Some Credit (GMC) [25] (financial insolvency prediction), and Lending Club [12](loan creditworthiness classification), see Appendix A for details. These datasets are selected for their diversity in dimensionality, feature types (mixed categorical and continuous variables), and real-world applicability, allowing us to evaluate the robustness and generalizability of our approach across different domains.
@@
For our experimental evaluation, we compare two variants of our method: DiCoFlex with default Euclidean distance and s-DiCoFlex which uses $L_{p}$ norm with $p=0.01$. This latter version encourages lower sparsity for continuous features, leading to explanations that alter fewer continuous features while maintaining validity, see Section 3.5.
@@
We compare our approach against several state-of-the-art counterfactual explanation methods. First of all, we consider methods, which generate multiple explanations for a single data point. CCHVAE [25] employs a conditional variational autoencoder to generate counterfactuals consistent with the data distribution. DiCE [19] uses gradient-based optimization to produce diverse counterfactuals in a model-agnostic framework. For all these methods, we generate up to 100 counterfactual explanations for each factual instance whenever feasible and randomly select 10 samples based solely on the validity criterion to provide an equal number of counterfactuals for every method.
Additionally, we compare with methods focused on returning a single explanation. ReViSE [13] introduces a probabilistic framework designed to separate sensitive and non-sensitive features with a focus on interpretability. The approach by Wachter et al. [40] represents a pioneering optimization-based method that minimizes a loss function balancing proximity to the input with the goal of changing the prediction. TABCF [22] specifically targets tabular data with a transformer-based VAE approach. This class of methods is expected to perform better because they are optimized for producing a single best explanation while the first class of methods takes a diversity criterion as a priority.
@@
We evaluate counterfactual quality using metrics aligned with key desirable properties. For test instances originally classified as class 0, we generate corresponding counterfactuals targeting class 1 and vice versa. Validity measures the percentage of counterfactuals that successfully change model predictions to the target class. Hypervolume [21] is used to measure diversity to quantify the spread and coverage of counterfactuals in the objective space. Plausibility is assessed through Local Outlier Factor (LOF) [3] that measure isolation from the training distribution. Sparsity quantifies feature modifications, calculated as the proportion of changed features in each category. For continuous features, we use $\varepsilon$-sparsity, which counts features modified beyond a threshold of 5% of the feature range. Proximity measures distance between original instances and the counterfactuals using the Euclidean distance only for continuous features. Observe that for categorical features proximity coincides with the sparsity. Probability measures model confidence in the target prediction. For the sparsity, proximity and plausibility metrics, lower values indicate better performance ($\downarrow$), while for the validity, probability and diversity metrics, higher values are preferable ($\uparrow$). We provide a detailed description of the metrics in Appendix B.

### v2

@@ 1 Introduction
Counterfactual explanations (CFs) have become an integral part of explainable artificial intelligence (XAI) by providing human-interpretable insights into complex machine learning models [guidotti2024counterfactual]. CFs answer critical "what-if" questions by suggesting minimal and meaningful changes to input data that could alter the outcome of a predictive model [Wachter2017counterfactual]. Such explanations have valuable applications across sensitive domains, including finance, healthcare, and legal decisions, where understanding the predictions of the model and potential alternative outcomes is paramount [zhou2023scgan, fontanella2024diffusion].
@@
According to guidotti2024counterfactual, an ideal counterfactual explanation should satisfy several key properties. First, it must demonstrate validity by successfully changing the model’s prediction. Second, it should maintain proximity by remaining as close as possible to the original input, minimizing the amount of change. Third, it should exhibit sparsity by modifying only the smallest possible set of features. Fourth, it must ensure plausibility by remaining within the realm of realistic, observed data. Finally, it should guarantee actionability by suggesting changes that are feasible for the stakeholder to implement.
In practice, however, these criteria often conflict. For instance, the smallest possible change may result in samples outside the realistic data distribution. To address this, mothilal2020explaining and laugel2023achieving argue that, rather than producing a single “best” counterfactual explanation, a diverse set of CFs should be generated. Presenting multiple realistic alternatives allows users to see different ways to achieve the desired outcome, explore trade‐offs between minimal change and realism, and gain deeper insight into the model’s decision boundaries. This diversity, in turn, empowers stakeholders to make more informed, feasible decisions.
Current approaches to diverse counterfactual generation face significant limitations. Optimization-based methods like DiCE [mothilal2020explaining] and multi-objective frameworks [dandl2020multi] incorporate diversity directly into their cost functions but require solving complex, separate optimizations for each explanation. This results in high computational costs and latency, making them impractical for real-time applications. Meanwhile, generative approaches such as those by pawelczyk2020learning and panagiotou2024TABCF offer efficiency improvements but frequently produce redundant explanations, require access to model gradients, and lack flexibility for adapting to user-defined constraints without extensive retraining.
@@ 2 Related Work
Counterfactual explanations (CFs) have become integral to explainable AI by providing intuitive alternatives that elucidate model decisions [Wachter2017counterfactual]. Optimization-based methods formulate CF generation as minimizing distance between instances while ensuring prediction changes, using implementations including gradient-based approaches [Wachter2017counterfactual], integer programming [ustun2019actionable], genetic algorithms [dandl2020multi], and satisfiability problems [karimi2020model]. Instance-based methods like FACE [poyiadzi2020face] select examples yielding desired predictions from existing data, while CBCE [keane2020good] merges features from paired contrasting instances. Model-specific approaches such as decision tree methods [guidotti2019factual, tolomei2017interpretable] leverage model structure by traversing alternative paths.
@@
Generating multiple diverse counterfactuals offers significant advantages over single explanations by providing alternative paths for recourse and better illustrating decision boundaries [laugel2023achieving, mothilal2020explaining]. DiCE [mothilal2020explaining] directly incorporates diversity into its loss function but requires separate optimization for each explanation, increasing computational costs. Other approaches achieve diversity through various strategies: DiVE [rodriguez2021beyond] employs determinantal point processes, while methods like [tsirtsis2020decisions, samoilescu2021model, carreira2021counterfactual] partition the feature space. Multi-objective approaches like MOC [dandl2020multi] generate solutions representing different trade-offs between properties such as proximity and sparsity, while CARE [rasouli2024care] and OrdCE [kanamori2021ordered] incorporate user preferences into the generation process.
@@
Generative models have gained popularity for counterfactual generation due to their ability to produce more plausible explanations that remain on the data manifold. VAE-based approaches include ReViSE [joshi2019towards], C-CHVAE [pawelczyk2020learning], CLUE [antoran2020getting], and CRUDS [downs2020cruds], which learn latent representations of the data distribution. GAN-based methods such as GeCo [schleich2021geco] and SCGAN [zhou2023scgan] use adversarial training to balance validity and realism. Recent transformer-based approaches like TABCF [panagiotou2024TABCF] specifically target tabular data with mixed types. Normalizing flows have also been explored for counterfactual generation [wielopolski2024probabilistically, furman2024unifying], where they model the class-conditional distribution for plausibility assessment during gradient-based optimization. However, many of these approaches still require access to model gradients, produce redundant explanations, or lack flexibility in handling user-defined constraints.
@@
Practical counterfactual explanations must satisfy various constraints including sparsity (minimizing modified features) and actionability (ensuring feasible changes). CERTIFAI [sharma2020certifai] employs genetic algorithms to enforce constraints, while MCCE [redelmeier2024mcce] conditions on immutable features. Methods focused on actionable recourse, such as CSCF [naumann2021consequence], account for downstream effects of feature modifications, while russell2019efficient and verma2021amortized generate explanations representing different modification strategies. However, most existing approaches require retraining or reoptimization when constraints change, limiting their flexibility in real-world applications.
@@
Current counterfactual methods face substantial limitations for tabular data: optimization-based approaches require high computational costs for diverse explanations [mothilal2020explaining], and generative models often produce redundant explanations, require model gradients, or lack constraint flexibility [pawelczyk2020learning]. Our proposed approach, DiCoFlex, addresses these limitations through a conditional generative framework that efficiently produces diverse counterfactuals in a single forward pass, enables real-time constraint customization, operates without requiring model access, and naturally handles mixed feature types.
@@ 3.2 Modeling predictive conditional distribution
Our approach aims at directly optimizing the conditional negative log-likelihood given by (4) with the data sampled from the distribution $q(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})$. As a consequence, it requires access to the parametrized density function. Several techniques, such as Kernel Density Estimation (KDE) and Gaussian Mixture Models (GMM), can be used to model the conditional density function $p_{\theta}(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})$. In this work, we propose using a conditional normalizing flow model [rezende2015variational] for this purpose. Unlike KDE or GMM, normalizing flows do not rely on a predefined parametric form of the density function and are well-suited for modeling high-dimensional data. Furthermore, in contrast to other generative models, such as diffusion models or GANs, normalizing flows allow for exact density evaluation via the change of variables formula and can be efficiently trained by minimizing the negative log-likelihood. The general formula that represents normalizing flows can be expressed as:
@@
where $f_{\theta}(\mathbf{x}^{\prime};\mathbf{x},y^{\prime})$ is an invertible transformation conditioned on $\mathbf{x}$ and $y^{\prime}$, $p_{Z}(f_{\theta}(\mathbf{x}^{\prime};\mathbf{x},y^{\prime}))$ is the base distribution, typically Gaussian and $J_{f_{\theta}}(\mathbf{x}^{\prime};\mathbf{x},y^{\prime})$ is Jacobian of the transformation. The biggest challenge in normalizing flows is the choice of the invertible function for which the determinant of Jacobian is easy to calculate. Several solutions have been proposed in the literature to address this issue with notable approaches, including NICE [DinhKB14], RealNVP [DinhSB17], and MAF [PapamakariosMP17]. Based on superior performance on benchmark dataset, we selected MAF as our normalizing flow model (see Appendix˜D).
3.3 Selecting training counterfactual explanations
The distribution $q(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime})$ is essential to sample data to train the flow model. Since we do not have direct access to this distribution, we introduce the straightforward technique of querying training counterfactual examples $\mathbf{x}^{\prime}$, which are similar to the input data and satisfy $h(\mathbf{x}^{\prime})=y^{\prime}$. For this purpose, we use a typical KNN (K-nearest neighbors) heuristic, in which we sample $K$ data points labeled as $y^{\prime}$ by the classifier that are closest to the input $\mathbf{x}$. Since we are sampling training examples from class $y^{\prime}$, the validity and plausibility constraints hold, while the proximity is satisfied by using the KNN approach.
@@ 3.3 Selecting training counterfactual explanations
$\hat{q}(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime},d)=\begin{cases*}\frac{1}{K}&if $\mathbf{x}^{\prime}\in N(\mathbf{x},y^{\prime},d,K)$,\\ 0&otherwise,\end{cases*}$ (6)
@@
One could ask whether a direct use of the distribution $\hat{q}(\mathbf{x}^{\prime}\mid\mathbf{x},y^{\prime},d)$ can be an alternative for generating counterfactual explanations, instead of learning a parametrized flow-based model $p_{\theta}(\mathbf{x}^{\prime}\mid\mathbf{x},y^{\prime})$}. Observe that the use of $\hat{q}$ is inherently limited by the number of available training examples and lacks the ability to generalize to unseen instances. Additionally, the diversity of generated examples is constrained by the number of selected neighbors. Modeling a generative model provides smooth changes over input examples and allows for extrapolating counterfactual generation outside training data. Importantly, our sampling strategy provides theoretical guarantees by construction: sampled counterfactuals are guaranteed to be valid, proximal, and plausible. We formalize these properties in Propositions A.3, A.4, and A.5 (Appendix A).
@@ 3.4 Sparsity on continuous features
[…]
We can observe that the level of modified attributes can be controlled through an appropriate choice of the value of $p$. For instance, if $p=0.01$, neighbors with fewer modified attributes are favored, thereby enforcing sparsity. Making use of $p=2$ gives a standard Euclidean proximity measure. Since categorical attributes are commonly encoded by one-hot vectors, the $L_{p}$ distance exhibits a discrete jump property: the distance equals 0 if categories match and a fixed value otherwise, regardless of $p$. This discrete structure naturally limits categorical changes. Therefore, the above scheme affects mostly continuous features.
In practice, the sparsity level can be effectively controlled in the inference time by incorporating the parameter $p$ into the flow model $p_{\theta}(\mathbf{x}^{\prime}\mid\mathbf{x},y^{\prime},p)$ through additional conditioning. During training, values of $p$ are sampled and explicitly included in the conditioning of the model, allowing it to learn the relationship between $p$ and the desired sparsity in the generated outputs. As a result, during the inference stage, users can adjust the sparsity level of the modified attributes simply by selecting an appropriate value of $p$. This mechanism provides a flexible and interpretable way to regulate the extent of changes applied to the input $\mathbf{x}$, depending on specific needs or constraints. Note, however, that the applied $L_{p}$ norm does not freeze the continuous attributes but only encourages small changes on their portion, which can be seen as a type of soft constraint. From a theoretical perspective, using small $p$ values (e.g., $p=0.01$) restricts the neighborhood $N(x,y^{\prime},d_{p},K)$ to a more concentrated region of the feature space, which typically corresponds to higher-density areas of the data manifold. This explains why DiCoFlex achieves improved plausibility scores empirically, as formalized in Proposition A.5.
@@ 3.6 Training
Concluding, the training procedure for estimating $\theta$ is as follows. For each training iteration, $\mathbf{x}$ is selected from data examples and the target class $y^{\prime}$ is sampled for counterfactual explanation. Value $p$ that represents the sparsity level is also sampled from some predefined set $\mathcal{P}$, as well as the masking vector $\mathbf{m}$ from $\mathcal{M}$. Next, the set $N(\mathbf{x},y^{\prime},d_{p,\mathbf{m}},K)$ of $K$ nearest neighbors is calculated considering distance measure $d_{p,\mathbf{m}}$ given by (8). The counterfactual $\mathbf{x}^{\prime}$ is sampled from $\hat{q}(\mathbf{x}^{\prime}|\mathbf{x},y^{\prime},d_{p,\mathbf{m}})$ given by (6). Finally, $\theta$ is updated in gradient-based procedure by optimizing (4), considering conditioning both on $p$ and m. Appendix˜E contains the details of the training algorithm.
The training procedure inherits several desirable theoretical properties from the $K$ nearest neighbors sampling mechanism. First, all training samples are guaranteed to be valid counterfactuals by construction (Proposition A.3). Second, proximity is controlled through an implicit upper bound determined by the $K$-th nearest neighbor (Proposition A.4). Third, plausibility is ensured as all samples originate from the training distribution (Proposition A.5). Finally, the normalizing flow learns to preserve diversity with a formal guarantee that expected diversity remains close to the empirical diversity, with deviation bounded by the training error (Theorem A.2). These theoretical foundations provide principled justification for the empirical performance demonstrated in Section 4.
@@ 4 Experiments
In this section, we conduct a series of experiments that evaluate the quality of counterfactual explanations generated by various methods [^1] Code available at https://github.com/ofurman/DiCoFlex. Our main objective is to understand how effectively these counterfactuals flip the predictions of a model while maintaining realism, requiring minimal changes to the input, and offering diversity among the generated alternatives.
@@ 4.1 Experimental Setup
We evaluate all methods on five well-established benchmark datasets commonly used in counterfactual explanation research: Adult [adult] (income prediction), Bank Marketing [moro2014data] (customer response classification), Default [default_of_credit_card_clients_350] (credit card default risk prediction), Give Me Some Credit (GMC) [pawelczyk2020learning] (financial insolvency prediction), and Lending Club [jagtiani2019roles] (loan creditworthiness classification), see Appendix˜B for details. These datasets are selected for their diversity in dimensionality, feature types (mixed categorical and continuous variables), and real-world applicability, allowing us to evaluate the robustness and generalizability of our approach across different domains.
@@
For our experimental evaluation, we compare two variants of our method: DiCoFlex with default Euclidean distance (p=2.0) and DiCoFlex (p=0.01), which uses the $L_{p}$ norm with p=0.01p{=}0.01. This latter version encourages higher sparsity for continuous features, leading to explanations that alter fewer continuous attributes while maintaining validity, see Section˜3.5. Importantly, these configurations are not separate methods but represent different inference-time settings of the same trained model. This design demonstrates DiCoFlex’s ability to dynamically adjust sparsity and proximity constraints without any retraining, enabling continuous user control over counterfactual behavior.
@@
We compare our approach against several state-of-the-art counterfactual explanation methods. First of all, we consider methods, which generate multiple explanations for a single data point. CCHVAE [pawelczyk2020learning] employs a conditional variational autoencoder to generate counterfactuals consistent with the data distribution. DiCE [mothilal2020explaining] uses gradient-based optimization to produce diverse counterfactuals in a model-agnostic framework. For all these methods, we generate up to 100 counterfactual explanations for each factual instance whenever feasible and randomly select 10 samples based solely on the validity criterion to provide an equal number of counterfactuals for every method.
Additionally, we compare with methods focused on returning a single explanation. ReViSE [joshi2019towards] introduces a probabilistic framework designed to separate sensitive and non-sensitive features with a focus on interpretability. The approach by Wachter et al. [Wachter2017counterfactual] represents a pioneering optimization-based method that minimizes a loss function balancing proximity to the input with the goal of changing the prediction. TABCF [panagiotou2024TABCF] specifically targets tabular data with a transformer-based VAE approach. This class of methods is expected to perform better because they are optimized for producing a single best explanation while the first class of methods takes a diversity criterion as a priority.
@@
We evaluate counterfactual quality using metrics aligned with key desirable properties. For test instances originally classified as class 0, we generate corresponding counterfactuals targeting class 1 and vice versa. Validity measures the percentage of counterfactuals that successfully change model predictions to the target class. Hypervolume [nowak2014empirical] is used to measure diversity to quantify the spread and coverage of counterfactuals in the objective space. Plausibility is assessed through Local Outlier Factor (LOF) [breunig2000lof] that measure isolation from the training distribution. Sparsity quantifies feature modifications, calculated as the proportion of changed features in each category. For continuous features, we use $\varepsilon$-sparsity, which counts features modified beyond a threshold of 5% of the feature range. Proximity measures distance between original instances and the counterfactuals using the Euclidean distance only for continuous features. Observe that for categorical features proximity coincides with the sparsity. Probability measures model confidence in the target prediction. For the sparsity, proximity and plausibility metrics, lower values indicate better performance ($\downarrow$), while for the validity, probability and diversity metrics, higher values are preferable ($\uparrow$). We provide a detailed description of the metrics in Appendix C.